An orange has 20 fewer calories than a banana. If 7 bananas have the same number of calories as 9 oranges, how many calories are in a banana?
Answer by priya(51)

FIRST form equations with the informations given in the question.

let O be the calories of orange and B be the calories of Banana.
eq 1. O = B - 20
eq 2. 7*B = 9* O
put the value of O from eq 1. in eq 2.,
we get,
7*B = 9*( B- 20)
7B = 9B - 180
REARRANGE THE VARIABLE ON ONE SIDE AND NUMBER ON THE OTHER SIDE.
we get,
180 = 9B -7B
180 = 2B
therefore B = 180/2
= 90 CALORIES IS THE ANSWER.
LEARN ONLINE - VISIT MY SITE
PRIYA